OpenStudy (anonymous):

3(x-3)-5x=-2(1+x)+x need a 2 column proof

OpenStudy (unklerhaukus):

LHS: \[3(x-3)-5x\]\[=3x-9-5x\]\[=-2x-9\] RHS: \[-2(1+x)+x \]\[=-2-2x+x\]\[=-2-x\] LHS≠RHS

OpenStudy (anonymous):

I thought the proof part was putting in the properties on the RHS?

OpenStudy (unklerhaukus):

i might have done this a different way to the way you want